OpenStudy (anonymous):

4 Algebra 2 problems... I give medals! :)

OpenStudy (anonymous):

1. Solve the system by substitution. 2x+y=-11 3x-4y=11

OpenStudy (anonymous):

2. Solve the system by elimination. 2x+6y=-12 5x-5y=10

OpenStudy (anonymous):

3. What is the solution of the following system? x-y=11 -x+y=-11

OpenStudy (anonymous):

4. What is the solution of the following system? -3x-2y=-12 9x+6y=-9

OpenStudy (anonymous):

The answer to each problem is either an ordered pair, no solution, or infinite solutions.

OpenStudy (gudden):

@Abbydavis102 this is the first question which is supposed to be solved by substitution so lets substitute the value of y from first equation into the second one 2x+y=-11 ............3x-4y=11 y = -11 - 2x substituting it in the second one we get 3x + 44 + 8x = 11 it means 11x = - 33 so u would get the value of x by solving it further and on equating this value in any of the above equation u would get the value of y too! :D

OpenStudy (gudden):

2. Solve the system by elimination. 2x+6y=-12..... 5x-5y=10 elimination is a method in which we are supposed to eliminate (remove/cancel) any of the one x or y from the equation so for the purpose multiply the first equation with 5 and the second one with 2 ..( this would make the number to be multiplied by x equal in both the cases.. and on subtracting the equations this would be eliminated) so the first equation comes out to be 10x + 30y = - 60 and the second one comes out to be 10x - 10y = 20 now on subtracting second equation from the first one we get 10x + 30y - (10x - 10y) = - 60 -(20) which comes out to be 40y = - 80 so solve it further to get the value of y.. and then equate it in any of the equations above to get the value of x..!

OpenStudy (gudden):

3. What is the solution of the following system? x-y=11 -x+y=-11 the answer of this question is that they have infinite solutions... because these equations are actually the same..!

OpenStudy (gudden):

4. What is the solution of the following s ystem? -3x-2y=-12 9x+6y=-9 from the second equation we can take 3 common on taking 3 common we get 3(3x +2y ) = 3 (-3) dividing both sides by 3 we get 3x + 2y = -3 from it we get the value of 2y = -3 - 3x and on equating it in the other equation ( -3x-2y=-12) u can get ur answer..!!

OpenStudy (gudden):

you would come to know if the equations have an ordered pair, no solution, or infinite solutions.

OpenStudy (anonymous):

Wow thanks!:)

OpenStudy (gudden):

welcome..:D